# A Closer Look into Mixture-of-Experts in Large Language Models

### v1

@@ 5.1 Outputs of Experts
Since the experts, ideally, are learned to specialize in different aspects, it is natural to question the similarities and differences between the outputs of selected and non-selected experts. In this experiment, we measure the correlation between the output feature vectors of experts using both input sequences. We plot the similarity heat maps for three tokens in the short sequence (Fig. 2) and the averaged heat map across all tokens in the long sequence (Fig. 3). Since the similarities have to be averaged for the long sequence, we employ angular similarity instead of cosine similarity for measurement so that the values range from 0 to 1:
@@
For observation convenience, the averaged similarity heat maps are plotted with the neuron-level similarity graphs of the expert weight matrices.
Mixtral and Mistral. The graphs of the short sequence show that the outputs from chosen experts tend to be more similar, which might be because their norms are generally larger. We will further discuss this in Sec 5.2. The overall similarities are relatively low in the deep (22nd-27th) layers, whereas many values significantly grow to greater than 0.8 in the last two layers. Furthermore, having multiple dark crosses in a graph is common, while the experts corresponding to the dark crosses can be more similar to the Mistral FFN (i.e., bright color in the last row). For the long sequence, the averaged heat maps have patterns similar to those of the neuron-level similarity graphs, including the dark crosses. The similarities also get lower with layer depth increasing, excepting the last layer. In addition, we have $S_{\text{ee}}>S_{\text{ef}}$ for both inputs. Most of the aforementioned observations are consistent with the previous analyses of static parameters (Sec 4.3), implying that measuring the similarity of weights, in some aspects, is equivalent to measuring the average similarity of outputs.
DeepSeek. Given the short input, most similarities are around zero, while the values in the last layer are significantly larger. Again, the similarities between experts chosen by the gate are likely to be higher, yet this difference emerges much less frequently than Mixtral. The averaged similarities of the long sequence also approach zero. Moreover, the amount of “small rectangular” with relatively light color in the graphs decreases when the layer gets deeper (except for the last layer), meaning that the averaged similarities are gradually lower.
Grok. Surprisingly, the similarities between the output features maintain a high level for all the tokens in the short sequence, indicating the experts are similar to each other in terms of behaviors. However, the similarities of their weight matrices are mostly zeros (Sec 4.1.1). We speculate that this is because the size of each Grok expert is relatively large. Each of them can learn comprehensive knowledge and behave alike despite distinct parameters. When averaging the similarities of the long input, some of the resulting averaged heat maps show similar patterns to the $W_{\text{gate}}$ figures. This relationship aligns with Mixtral’s observation.
@@ 5.2 Norms of Expert Outputs and Gate Scores
Figure 4: The experts’ L2 norms and the gate scores of the short sequence. Each token’s top $k$ experts are shown on top of each heat map. Each number in the horizontal axis refers to an expert index.
@@
In Sec 5.1, we find that the outputs from chosen experts tend to be more alike. To investigate possible reasons for this observation, we study the relationship between the experts’ L2 norm and the gate decision in this experiment. We employ the short sequence to be the input. The calculated norms, along with the gate scores, are plotted in Fig. 4.
@@
Mixtral. We found that the two experts chosen by the gate usually output feature vectors with the highest norms, which reveals that the norm might be one of the key factors for gate decisions. This finding agrees with the router’s design in CompeteSMoE Pham et al. (2024), which selects experts based on their output norms. It can also be regarded as an explanation for why the outputs of the chosen Mixtral and DeepSeek experts tend to be more alike (Sec 5.1). In Append D, we repeat this experiment using the long input, and the statistical results further demonstrate the “higher norm, higher score” observation. In Fig. 4, we can also see that the gate scores assigned for the top-1 experts are usually much higher than the others, including the second place. This demonstrates that the gate is learned to strengthen the confidence of decision during the training process. On the other hand, the deeper the layer, the larger the norm, which is similar to the growth in standard models Shleifer et al. (2021).
@@
DeepSeek. In contrast to the observation about Mixtral experts, the gate decision seems to depend less obviously on the output norms of DeepSeek experts. However, the top-1 experts often score much higher than the remaining candidates. The magnitude of norms is proportional to the depth, yet the increment is smaller than Mixtral. In the last layer, the variance of norms becomes greater.
@@
Grok. While the scores of the Top-1 experts are higher than the others, no correspondence between the norms and the gate scores is observed. One of the possible reasons might be the relatively low activation ratios of GeLU (see Sec 5.3) result in weaker dependence on the norm for the gate decisions. Besides, unlike Mixtral and DeepSeek, the magnitude of the norms hardly changes across the depth, and some of the norm values can be smaller than 1, which is rare in the other two models.
5.3 Intermediate States of Experts
While Sec 5.1 has studied the final outputs of experts, we continue to analyze their intermediate outputs here to examine the inner states of experts. Given an input $x$, the intermediate state of an expert refers to the output of $\operatorname{act}(W_{\text{gate}}x)$, which is a $d_{\text{hid}}$ dimensional vector. These intermediate vectors control the activation of neurons so we simply record them for analysis. The short sequence is used as the input. Mixtral, Mistral, and DeepSeek utilize SiLU as the activation function while Grok adopts GeLU. Fig. 5 depicts the magnitude of the vectors. Due to the page limit, we only plot the results of Mixtral for three tokens.
Common. There are some horizontal lines in each figure, which means there is an “out-liner” expert with the highest or lowest activation values. Nonetheless, there is no clear relation between such phenomena and gate decisions.
Mixtral and Mistral. For a single token, we found that the absolute activation value of 99.6% elements in each expert are greater than 0.001 after SiLU activation function on average. Such large ratio indicates that the vast majority of neurons in an expert are activated. In Fig. 5, some vertical lines across all experts are commonly found, meaning that the $W_{\text{gate}}$ matrices of different experts assign similar activation values to neurons with the same indices. In addition, the magnitude of the intermediate states grows along with the layer depth, similar to the observation in Sec 5.1.
DeepSeek. On average, each expert’s absolute activation value of 99.7% neurons reaches beyond 0.001 after SiLU. Vertical lines rarely exist in the DeepSeek model. Similarly, the elements in the intermediate state vectors get larger as the layer goes deeper.
Grok. Employing GeLU as the activation function, only 25.3% neurons per Grok expert attain an absolute activation value greater than 0.001. The activation values are generally smaller than Mixtral and DeepSeek. Li et al. (2022); Song et al. (2024a) suggest such difference largely comes from different activation functions. Interestingly, Song et al. (2024b) further utilize the sparsity in experts in SMoE and achieve SOTA performance when activate the same number of parameters.
@@ 5.4 Summary
The observations of dynamic behaviours are concluded below: (1) The outputs of Mixtral and DeepSeek experts in deep/last layers are less/much alike. This can be observed from the heat maps for both the short (Fig. 2) and long (Fig. 3) inputs. (2) The averaged heat maps of expert outputs have similar patterns to the neuron-level similarity graphs (Fig. 3), implying that measuring the similarity of weights, in some aspects, is equivalent to measuring the average similarity of outputs. (3) The outputs of Grok experts are highly similar (Fig. 2), which may be due to their large sizes. (4) For both Mixtral and DeepSeek, experts that output feature vectors with larger norms are likely to achieve higher gate scores, as shown in Fig. 4. We further verified this observation in Fig. 9 in the Appendix. (5) For Mixtral, neurons of different experts positioned at the same indices have similar activation values. In Fig. 5, some vertical lines can be found in the Mixtral heat maps.
@@ 6 Discussion
Based on our analyses, we provide some suggestions for MoE models in various aspects:
  • Neuron-level experts: Intuitively, the gate embedding determines the expert selection while $W_{\text{gate}}$ is responsible for choosing neurons to activate. Meanwhile, we found that the similarities of the gate embedding and $W_{\text{gate}}$ show association. This implies that the neurons might be some more fine-grained experts. Therefore, operations on experts, such as division, construction, and composition, should be further studied at the micro level.
  • Model architecture: Since the similarities between experts tend to be relatively low/high in deep/last layers, one can consider increasing the number of experts in deep layers while reducing it in the last layers. In addition, the gate is found to select experts with larger output norms frequently, so employing the norm as the routing mechanism is reasonable (Pham et al. (2024) have empirically proven the effectiveness).
  • Correlation measurement: When analyzing the correlation between experts, measuring the similarities between their weight matrices gives partly equivalent results to measuring the similarities of their output feature vectors over considerable tokens. Hence, measuring the weight matrices can obtain an overview while inspecting the outputs of various tokens individually, which benefits fine-grained investigation.
  • Training scheme: The Mixtral training method has not been publicly announced. However, we observed some attributes shared by Mixtral experts (e.g., relatively high similarities of weight matrices), and certain relationship between experts and the FFN of Mistral (e.g., similar intermediate states). Therefore, we conjecture that the Mixtral model is trained by some special initialization schemes other than from scratch, e.g., upcycling Komatsuzaki et al. (2022) from Mistral, that is, copying all experts from the FFN. On the contrary, the experts of DeepSeek and Grok, who are known to be trained from scratch, show a weaker correlation than Mixtral experts in our experiments. Hence, we speculate that compared with some initialization approaches, training a MoE model from scratch shows stronger potential to facilitate the diversification of experts.
@@ 7 Related Work
Most existing works analyze MoE from the router’s perspective by observing expert selections. Early works have observed the unstable choices in the router Zuo et al. (2021); Chi et al. (2022); Dai et al. (2022). Recent works find the standard router does not show clear specialization in the domain level Jiang et al. (2024); Dai et al. (2024) and mainly route based on token ID instead of high-level semantics Xue et al. (2024). Some works investigate the expert’s similarity Wu et al. (2022), discovering and utilizing redundancies in experts for efficient inference Li et al. (2023); Lu et al. (2024). Liu et al. (2023); Qiu et al. (2023) notice the connection between routing connection and expert computation, and utilize the average of the experts’ first layer weights to guide routing. Pham et al. (2024) proposes adding the expert’s output norm as a supervision signal for routing training. These works provide insights into MoE from one or two viewpoints, while our work offers a systematic analysis and comparison.
@@ 8 Conclusion
In this paper, we initially attempt to investigate the inner working mechanisms of MoEs by studying the parameters and outputs of three different MoE models. We conclude the empirical observations and propose some practical suggestions for various aspects. It is still too early to answer whether MoEs do learn heterogeneous experts. Yet some of our experiments suggest that specific architecture designs (e.g., the number of experts) and training frameworks might be more beneficial to expert specialization. We hope this work can provide some inspiring insights and serve as a valuable foundation for future research about MoE and other modular architectures.
@@ 9 Limitations
Our analysis is mainly based on observation and lacks intervention for further validation.
Appendix A Projection of Expert Matrices in Low-dimensional Space
A.1 Matrix-level
@@
To better understand the expert relationship, we employ principal components analysis (PCA) to convert the flattened vectors of weight matrices to two-dimensional space. The vectors are standardized before applying PCA. Fig. 6 depicts the projection in 2D plane.
Mixtral and Mistral. Similar to the observation in Sec 4.1.1, the figures of the three matrices are alike. In general, about half of the Mixtral experts stay close to each other and to Mistral FFN, while the others locate much farther away. Moreover, the outliers correspond with the dark crosses.
DeepSeek. Only routed experts are considered due to different hidden sizes. Since several outliers exist and cause the remaining data points to be densely gathered, we remove them using the DBSCAN algorithm with $\epsilon=50$ and plot the rest in Fig. 6. It can be observed that the experts distribute rather densely, especially for $W_{\text{up}}$. The distribution of experts varies for three matrices, but the figures of $W_{\text{up}}$ and $W_{\text{down}}$ are more alike than those of the gate matrix.
Grok. Typically, about half of the Grok experts densely gather for $W_{\text{up}}$ and $W_{\text{down}}$. Another half turns out to be outliers although no dark cross is observed before. Furthermore, the outliers of the three matrices are partly coincided.
A.2 Neuron-level
@@
To project the neurons onto a 2D/3D space, each row vector of $W_{\text{up}}$ and $W_{\text{gate}}$, or column vector of $W_{\text{down}}$, is treated as a single data point. Standardization is then applied, following by the PCA. Visualization of the principle components is illustrated in Fig. 7. Different colors refer to neurons belong to different experts.
Common. The vast majority of neurons gather in the low dimensional space. In some of the layers, the distribution of neurons forms a special shape, such as a cross or a thick line, which appears the most often for $W_{\text{down}}$, followed by $W_{\text{up}}$ and finally $W_{\text{gate}}$. Comparing with ellipse, these shapes indicate that the neurons are relatively more similar.
Mixtral and Mistral. The neurons of Mistral FFN distribute more densely than those of Mixtral experts. Notably, the distribution shape of neurons in FFN and experts are usually alike, even for the outliers.
DeepSeek and Grok. The number of outliers is a bit greater than Mixtral.
Appendix B Averaging Expert Neurons
To investigate the expert correlation in neuron level, the averaging approach simply averages the rows (for $W_{\text{up}}$ and $W_{\text{gate}}$) or the columns (for $W_{\text{down}}$) of the weight matrices and then calculates the similarity of the resulting vectors across experts. Fig. 8 displays the graphs.
Common. The heat maps of $W_{\text{up}}$ and $W_{\text{down}}$ are almost the same as in Section 4.1.1. Yet the similarities of $W_{\text{gate}}$ significantly increase.
Mixtral and Mistral. The dark crosses sometimes disappear. In the figures of $W_{\text{gate}}$, the similarities between the experts and the Mistral FFN are often lower than the ones between the experts (i.e., $S_{\text{ee}}>S_{\text{ef}}$), which is contrary to the previous observation. This can happen if the expert neurons in different positions are alike. For instance, given three vectors $f=(0,0)$, $e_{1}=(1,0)$, and $e_{2}=(0,1)$, the vector similarity $S_{e_{1}e_{2}}$ is lower than $S_{e_{1}f}$ and $S_{e_{2}f}$. If averaging the elements, we have $\bar{f}=(0)$, $\bar{e}_{1}=(0.5)$, and $\bar{e}_{2}=(0.5)$, then $S_{e_{1}e_{2}}$ becomes the highest.
DeepSeek. The growth of the $W_{\text{gate}}$ similarity values is directly proportional to the depth of layer.

### v2

@@ 5.1 Outputs of Experts
Since experts are ideally learned to specialize in different aspects, it is natural to question the similarities and differences between the outputs of selected and non-selected experts. In this experiment, we measure the correlation between the output feature vectors of experts. We plot the similarity heat maps for three tokens in the short input (Fig. 2) and the average heat map across all tokens in the long input (Fig. 3). For the long input, we use angular similarity instead of cosine similarity for measurement, as the similarities need to be averaged, ensuring that the values range from 0 to 1:
@@
For clarity, the average similarity heat maps are plotted alongside the neuron-level similarity graphs of the expert weight matrices. Fig. 9 further depicts the results from additional datasets, which are consistent with those of the long input.
[…]
DeepSeek. Given the short input, most similarities are around zero, while the values in the last layer are significantly larger. Again, the similarities between experts chosen by the gate are likely to be higher, although this difference occurs much less frequently than in Mixtral. The average similarities for the long input also approach zero. Moreover, the number of “small rectangular” with relatively light color in the graphs decreases as the layer depth increases (except for the last layer), meaning that the average similarities gradually decline.
Grok. Surprisingly, the similarities between the output features remain high for all tokens in the short input, indicating the experts exhibit similar behaviours. However, the similarities of their weight matrices are mostly zeros (§ 4.1.1). We speculate that this may be due to the relatively large size of each Grok expert, allowing each to learn comprehensive knowledge and behave similarly despite having distinct parameters. When averaging the similarities for the long input, some of the resulting average heat maps display patterns similar to those of the $W_{\text{gate}}$ figures. This relationship aligns with the observations made for Mixtral.
@@ 5.2 Norms of Expert Outputs and Gate Scores
Figure 4: The experts’ L2 norms and the gate scores of the short input. Each token’s $k$ experts are shown on top of each heat map. Each number in the horizontal axis refers to an expert index.
@@
In § 5.1, we find that the outputs from chosen experts tend to be more alike. To investigate the possible reasons for this observation, we employ the short input to study the relationship between the experts’ L2 norm and the gate decision in this experiment. The calculated norms, along with the gate scores, are plotted in Fig. 4. In Append G, we repeat this experiment using the long input and additional datasets, and the results also support the “higher norm, higher score” observation.
@@
Mixtral. We found that the two experts chosen by the gate usually output feature vectors with the highest norms, which reveals that the norm might be one of the key factors in gate decisions. This finding agrees with the router’s design in CompeteSMoE Pham et al. (2024), which selects experts based on their output norms. It also helps explain why the outputs of the chosen Mixtral and DeepSeek experts tend to be more alike (§ 5.1). In Fig. 4, we observe that the gate scores assigned to the top-1 experts are usually much higher than those of the others, including the second place. This demonstrates that the gate is learned to strengthen the confidence of its decision during the training process. On the other hand, the deeper the layer, the larger the norm, which is similar to the growth in standard models Shleifer et al. (2021).
@@
DeepSeek. In contrast to the observation about Mixtral experts, the gate decision appears to depend less obviously on the output norms of DeepSeek experts. However, the top-1 experts often score much higher than the remaining candidates. The magnitude of the norms increases with depth, although the increment is less pronounced than in Mixtral. In the last layer, the variance of norms becomes greater.
@@
Grok. While the scores of the top-1 experts are higher than those of the others, no correspondence between the norms and the gate scores is observed. One possible reason could be the relatively low activation ratios of GeLU (see Append H), which may lead to a weaker dependence on the norm for gate decisions. Besides, unlike Mixtral and DeepSeek, the magnitude of the norms hardly changes across depth, and some of the norm values can be less than 1, which is rare in the other two models.
Figure 5: Normalized model performance across benchmarks for the dynamic expert numbers experiment. Solid line represents lower is better, while dashed line represents higher is better. “Bench avg” refers to the average performance over the four benchmarks evaluated.
@@ 5.3 Summary
The observations of dynamic behaviours are concluded below: i) The outputs of Mixtral and DeepSeek experts in deep (last) layers are less (much) alike. This can be seen in the heat maps for both the short (Fig. 2) and long (Fig. 3) inputs. ii) The average heat maps of expert outputs resemble the neuron-level similarity graphs (Fig. 3), implying that weight similarity measurements can reflect output similarity. iii) Grok experts exhibit high output similarity (Fig.2), likely due to their larger sizes. iv) For Mixtral and DeepSeek, experts generating feature vectors with larger norms tend to receive higher gate scores, as shown in Fig. 4. We further verified this observation in Fig. 10.
@@ 6 Discussion
Based on our analyses, we offer several suggestions for MoE models across various aspects.
Neuron-level experts. Intuitively, the gate embedding determines expert selection while $W_{\text{gate}}$ is responsible for choosing which neurons to activate. Meanwhile, we found that the similarities of the gate embedding and $W_{\text{gate}}$ show association. This implies that neurons may function as more fine-grained experts. Therefore, operations on experts, such as division, construction, and composition, should be further studied at the micro level. For instance, MoEfication Zhang et al. (2021) and EMoE Qiu et al. (2024a) construct MoE experts by splitting the MLP layers of a dense model, suggesting our findings from a similar perspective.
Model architecture. Given that the similarities between experts tend to be relatively low (high) in deep (last) layers, one can consider increasing the number of experts in the deeper layers while reducing it in the last layers. In addition, since the gate frequently selects experts with larger output norms, employing norm-based routing mechanism is a reasonable approach. Empirical evidence from Pham et al. (2024) supports this effectiveness.
[…]
Correlation measurement. Analyzing expert correlations through weight matrix similarities yields partially equivalent results to those from output feature vector similarities across considerable tokens. Thus, assessing weight matrices offers a broader overview, while examining individual token outputs allows for more detailed analysis.
Training scheme. The training method for Mixtral has not been publicly announced. However, we observed certain characteristicss shared by Mixtral experts (e.g., relatively high similarities of weight matrices), and a notable relationship between these experts and the Mistral FFN (e.g., similar intermediate states in Fig. 12). Consequently, we conjecture that the Mixtral model may be trained using special initialization schemes other than from scratch, e.g., upcycling Komatsuzaki et al. (2022) from Mistral, that is, copying all experts from the FFN. On the contrary, the experts of DeepSeek and Grok, which are known to be trained from scratch, show weaker correlations than Mixtral experts in our experiments. Similarly, Wei et al. (2024a) tracks changes in expert similarities throughout the training process, observing that upcycled experts exhibit greater similarity compared to those randomly initialized. Hence, we speculate that training a MoE model from scratch shows stronger potential to facilitate the diversification of experts compared with certain initialization approaches.
@@ 7 Related Work
Due to the page limit, we focus on existing works analyzing MoEs. An extended related work section for MoE LLMs can be found in Append B.
Most existing works analyze MoE from the router’s perspective by observing expert selections. Early works have observed the unstable choices in the router Zuo et al. (2021); Chi et al. (2022); Dai et al. (2022). More recent studies find the standard routers do not show clear specialization at the domain level Jiang et al. (2024); Dai et al. (2024) and primarily route based on token ID instead of high-level semantics Xue et al. (2024). Shi et al. (2024) shows that Top-2 and Rank-k routing result in different model behaviours and propose a new self-contrast decoding method to determine the next-token distribution based on this finding.
Other works investigate the expert’s similarity Wu et al. (2022), uncovering and utilizing redundancies among experts for efficient inference Li et al. (2023); Lu et al. (2024). Zhang et al. (2024) reveals the redundancy within experts and perform pruning based on their similarities. Liu et al. (2023); Qiu et al. (2023) notice the connection between routing connection and expert computation, and utilize the average of the experts’ first-layer weights to guide routing. Pham et al. (2024) proposes adding the expert’s output norm as a supervision signal for routing training.  Chen et al. (2022) empirically and theoretically proves that a two-layer MoE CNN is able to learn cluster-center features via specializing experts to specific portions of the data. While these works provide insights into MoE from one or two viewpoints, our work offers a systematic analysis and comparison focusing on transformer-based MoE LLMs.
@@ 8 Conclusion
In this paper, we initially attempt to investigate the inner working mechanisms of MoEs by studying the parameters and outputs of three different MoE models. We summarize our empirical observations and propose practical suggestions across various aspects. While it is premature to conclude whether MoEs genuinely learn heterogeneous experts, some of our experiments indicate that specific architectural designs (e.g., the number of experts) and training frameworks may facilitate expert specialization. We hope this work can provide inspiring insights and serve as a valuable foundation for future research on MoE and other modular architectures.
@@ 9 Limitations
The limitations of our work include: 1) Although the models we investigated cover several common designs of MoE, our analysis does not encompass all aspects (e.g., other routing strategies like top-1 routing or model architectures that place MoE layers at every other layer); 2) Despite the availability of other metrics, we primarily adopt cosine similarity in our experiments involving similarity measurement, as it is a widely used approach Pham et al. (2024); Chen et al. (2022); 3) We mainly focus on the pretrained base model but seldom explore the behaviours of models after fine-tuning. Analyzing the changes in expert behaviours during the fine-tuning process could yield valuable insights.
Appendix A Model Selection
Our experiments are conducted on Mixtral 8x7B, DeepSeekMoE, and Grok-1. We choose these models due to their widespread use and impressive performance across various domains. Additionally, these models are complementary in several crucial attributes, such as training scheme, activation functions, top-k settings, and the number of experts, as listed in Tab 1 and Tab 4. This allows for a comparative analysis with controlled variables and encompasses a wide range of parameter sizes, from rather small (16B) to relatively huge (314B). Hence, we believe that the findings derived from these three models are fairly robust, despite the limited number of models examined.
Model Training scheme Activation # Total layers # Total params # Activated params
Mixtral unknown (upcycling) SiLU 32 46.7B 12.9B
Mistral from scratch SiLU 32 7.3B 7.3B
DeepSeek from scratch SiLU 28 16.4B 0.3B
Grok from scratch GeLU 64 314B 78.5B
Table 4: Additional information of chosen models.
Appendix B Extended Related Work
MoE LLMs. MoEs have garnered significant attention in recent years due to their ability to efficiently scale model capacity with minimal computational overhead. Most current transformer-based MoE LLMs adopt a typical architecture design that replaces the original FFN with multiple expert networks and a sparse gating network Wei et al. (2024a); Wu et al. (2024); Dai et al. (2024); Xue et al. (2024); Jiang et al. (2024); Zoph et al. (2022). JetMoE Shen et al. (2024) and ModuleFormer Shen et al. (2023) incorporate Mixture of Attention Heads Zhang et al. (2022) into their model, achieving further sparsity. A recent survey Cai et al. (2024) provides a comprehensive review of both the algorithmic and system design aspects of MoEs. For this study, we select three representative candidates among current open-sourced MoE LLMs for analysis to gain intriguing insights.
Appendix C Projection of Expert Matrices in Low-dimensional Space
C.1 Matrix-level
@@
To better understand the relationships among experts, we employ principal components analysis (PCA) to project the flattened vectors of weight matrices into two-dimensional space. The vectors are standardized before applying PCA. Fig. 6 depicts the resulting 2D projection.
Mixtral and Mistral. Consistent with the observations in § 4.1.1, the figures for the three matrices appear similar. Generally, about half of the Mixtral experts cluster closely together and near the Mistral FFN, while the others locate much farther away. Moreover, the outliers correspond to the dark crosses.
DeepSeek. Only routed experts are considered due to differences in hidden sizes. Because several outliers exist, causing the remaining data points to be densely gathered, we remove them using the DBSCAN algorithm with $\epsilon=50$ and plot the rest in Fig. 6. It can be observed that the experts distribute rather densely, especially for $W_{\text{up}}$. Although the distribution of experts varies for three matrices, the figures for $W_{\text{up}}$ and $W_{\text{down}}$ are more similar than those of the gate matrix.
Grok. Typically, about half of the Grok experts densely gather for $W_{\text{up}}$ and $W_{\text{down}}$. The other half turns out to be outliers even though no dark cross were observed before. . Furthermore, the outliers of the three matrices partially coincide.
C.2 Neuron-level
@@
To project the neurons into a 2D or 3D space, each row vector of $W_{\text{up}}$ and $W_{\text{gate}}$, or each column vector of $W_{\text{down}}$, is treated as a single data point. Standardization is then applied, following by PCA. The visualization of the principal components is illustrated in Fig. 7. Different colors refer to neurons belonging to different experts.
Common. The vast majority of neurons gather in the low-dimensional space. In some layers, the distribution of neurons forms a special shape, such as a cross or a thick line, which appears the most often for $W_{\text{down}}$, followed by $W_{\text{up}}$, and finally $W_{\text{gate}}$. Compared to ellipses, these shapes indicate that the neurons are relatively more similar.
Mixtral and Mistral. The neurons in the Mistral FFN distribute more densely than those of the Mixtral experts. Notably, the distribution shape of neurons in the FFN and experts are usually alike, even for the outliers.
DeepSeek and Grok. The number of outliers is a bit greater tahn that observed in Mixtral.
Appendix D Averaging Expert Neurons
To investigate expert correlation at the neuron level, the averaging approach simply averages the rows (for $W_{\text{up}}$ and $W_{\text{gate}}$) or the columns (for $W_{\text{down}}$) of the weight matrices and then calculates the similarity of the resulting vectors across experts. Fig. 8 displays the graphs.
Common. The heat maps of $W_{\text{up}}$ and $W_{\text{down}}$ are nearly identical to those presented in § 4.1.1. Yet the similarities of $W_{\text{gate}}$ significantly increase.
Mixtral and Mistral. The dark crosses sometimes disappear. In the figures for $W_{\text{gate}}$, the similarities between the experts and the Mistral FFN are often lower than the similarities among the experts themselves (i.e., $S_{\text{ee}}>S_{\text{ef}}$), which is contrary to previous observations. This can happen if the expert neurons in different positions are alike. For instance, given three vectors $f=(0,0)$, $e_{1}=(1,0)$, and $e_{2}=(0,1)$, the vector similarity $S_{e_{1}e_{2}}$ is lower than $S_{e_{1}f}$ and $S_{e_{2}f}$. If averaging the elements, we have $\bar{f}=(0)$, $\bar{e}_{1}=(0.5)$, and $\bar{e}_{2}=(0.5)$, then $S_{e_{1}e_{2}}$ becomes the highest.
DeepSeek. The growth of the $W_{\text{gate}}$ similarity values is directly proportional to the layer depth.